OpenStudy (anonymous):

X-4. 4. Solve with all work, ______ = ___ 3. 6.

jimthompson5910 (jim_thompson5910):

Start off by cross multiplying to get 6(x-4) = 3*4 6x-24 = 12 what's next?

OpenStudy (anonymous):

I'm really bad with algebra. I'm not sure what it is your wanting me to multiply.

jimthompson5910 (jim_thompson5910):

we want to isolate x so we want x all by itself

jimthompson5910 (jim_thompson5910):

how do we do this

OpenStudy (anonymous):

So I'd subtract all the numbers around it ??

jimthompson5910 (jim_thompson5910):

kinda more like add 24 to both sides (to undo the -24) what do you get when you do this?

OpenStudy (anonymous):

By adding 2. If we add 24 tonegitive 24.... It would be zero

jimthompson5910 (jim_thompson5910):

no, we add 24 to both sides to get 6x-24 = 12 6x-24+24 = 12+24 6x = 36 what's next?

OpenStudy (anonymous):

Subtracting by the left side?

OpenStudy (anonymous):

I meant right side.

jimthompson5910 (jim_thompson5910):

how do you undo multiplication?

OpenStudy (anonymous):

Division.

jimthompson5910 (jim_thompson5910):

so you divide both sides by 6

jimthompson5910 (jim_thompson5910):

6x = 36 x = 36/6 x = 6

jimthompson5910 (jim_thompson5910):

x is finally isolated and all by itself, so you're done